OpenStudy (anonymous):

Find the slope of the line that contains the points named. (Use fraction format.) A(3, 2), B(7, 8) m =

OpenStudy (amistre64):

A( 3, 2) -B(-7,-8) ---------- -4,-6; m=6/4 = 3/2